OpenStudy (anonymous):

how can substitution be used to solve a system of equations that doesn't have a coefficient of 1 or -1

OpenStudy (anonymous):

Is this the homework question or your interpretation of it?

OpenStudy (anonymous):

I could explain how substitution works...

OpenStudy (anonymous):

i know how it works i just don't get the question

OpenStudy (anonymous):

Solving a single linear equation is getting a variable by itself on one side of an equation. An example of a one-step linear equation is x+3=5. Subtracting 3 from both sides of the equation gets x by itself on the left side, and results in 2 on the right side. Combining like terms may be necessary to get x by itself. For example, x+2+x+6=14. Combining the x's with each other, and the constant numbers with each other changes the equation to: 2x+8=14. This is a two-step equation, because the x is being multiplied by a number, as well as a constant being added. Subtracting 8 gives 2x=6. Dividing by 2 gives x=3

OpenStudy (anonymous):

I don't either. You just do substitution. Solve for one variable in one equation, and plug that into the other equation...

OpenStudy (anonymous):

he first step you should do is to simplify both sides of the equation by collecting the common terms containing variable x and the common constant terms separately at each side of the equation. The second step is to collect terms containing variable x in one side of the equation and make the other side of the equation free of variable terms. The third step is to collect all the constant terms in the right side and make the left side free of constants. The fourth step is to divide both sides of your equation by the coefficient at x. You will get an expression of the form x = the value, which is going to be a solution. The last step in solving the equation is to check the solution. In order to do it, simply substitute the found value of the variable x into the original equation, then calculate and compare both sides.

OpenStudy (anonymous):

thx both of u

OpenStudy (anonymous):

cut and paste job lol

OpenStudy (anonymous):

These are equations with just a plain variable like 'x', rather than something harder like 'x^2'. They are the simplest forms you'll deal with. For example '--- + 2=5'.

OpenStudy (anonymous):

@david111

OpenStudy (anonymous):

Solving linear equations is applying a particular solving technique in order to find out the "x" and "y" in the linear equation. This can be done by plotting on the graph. When there are more than one corresponding linear equations, for e.g Ax + By + C = 0 - Eq1 and Dx + Ey + F = 0 - Eq2 In such a case we can substitute the rearranged value of "x" from Eq1 into Eq2 or vice-versa. This will help in calculating the values of "x" and "y".

OpenStudy (anonymous):

In applications of math it is rare for the answer to come out neatly, by being able to solve linear equations you can take equations and variables and find the value of each variable.

OpenStudy (anonymous):

can u dumb that down a little lol

OpenStudy (anonymous):

The idea here is to solve one of the equations for one of the variables, and plug this into the other equation. It does not matter which equation or which variable you pick. There is no right or wrong choice; the answer will be the same, regardless. But — some choices may be better than others. For instance, in this case, can you see that it would probably be simplest to solve the second equation for "y =", since there is already a y floating around loose in the middle there? I could solve the first equation for either variable, but I'd get fractions, and solving the second equation for x would also give me fractions. It wouldn't be "wrong" to make a different choice, but it would probably be more difficult. Being lazy, I'll solve the second equation for y

OpenStudy (anonymous):

theres no problem with being lazy

OpenStudy (anonymous):

http://www.sosmath.com/soe/SE/SE.html